# for Generative QA

### v1

@@ 6 Conclusions
We proposed $\beta$-calibration, a new notion of calibration which conditions on groups of QA pairs. We propose two new posthoc calibration schemes for LM-elicited confidence scores. Our algorithms are effective on various QA datasets. For future work, we plan to investigate alternative notions of calibration for groups, such as multi-calibration. Other choices of $\beta$ that generalize expected calibration error can also be used, such as random projection tree (Dasgupta and Freund, 2008), which adapts to the intrinsic low-dimensional structure in the data.
@@
Dataset Prompt LLM Calibrator $\text{CE}(h;\beta)$ AUAC MMLU Ling1s-Top1 Mistral BB (ours) $0.171\pm 0.005$ $0.20\pm 0.023$ MMLU Ling1s-Top1 Mistral HS-BB (ours) $\mathbf{0.16\pm 0.005}$ $\mathbf{0.269\pm 0.022}$ MMLU Ling1s-Top1 Mistral S-BB (ours) $0.163\pm 0.003$ $0.19\pm 0.045$ MMLU Ling1s-Top1 Mistral S-B $0.393\pm 0.004$ $0.141\pm 0.003$ MMLU Ling1s-Top1 Mistral S $0.249\pm 0.005$ $0.122\pm 0.002$ MMLU Ling1s-Top1 Mistral B $0.392\pm 0.003$ $0.139\pm 0.007$ MMLU Ling1s-Top1 Mistral None $0.532\pm 0.008$ $\mathbf{0.269\pm 0.007}$ MMLU Ling1s-Top1 Gemma BB (ours) $0.232\pm 0.005$ $0.211\pm 0.032$ MMLU Ling1s-Top1 Gemma HS-BB (ours) $0.182\pm 0.005$ $0.26\pm 0.01$ MMLU Ling1s-Top1 Gemma S-BB (ours) $\mathbf{0.181\pm 0.009}$ $\mathbf{0.275\pm 0.029}$ MMLU Ling1s-Top1 Gemma S-B $0.382\pm 0.006$ $0.184\pm 0.002$ MMLU Ling1s-Top1 Gemma S $0.201\pm 0.009$ $0.19\pm 0.003$ MMLU Ling1s-Top1 Gemma B $0.385\pm 0.003$ $0.183\pm 0.008$ MMLU Ling1s-Top1 Gemma None $0.603\pm 0.006$ $0.249\pm 0.01$ MMLU Verb1s-Top1 Mistral BB (ours) $0.197\pm 0.004$ $0.198\pm 0.046$ MMLU Verb1s-Top1 Mistral HS-BB (ours) $\mathbf{0.149\pm 0.004}$ $\mathbf{0.306\pm 0.035}$ MMLU Verb1s-Top1 Mistral S-BB (ours) $0.16\pm 0.004$ $0.217\pm 0.057$ MMLU Verb1s-Top1 Mistral S-B $0.362\pm 0.005$ $0.139\pm 0.004$ MMLU Verb1s-Top1 Mistral S $0.258\pm 0.005$ $0.126\pm 0.001$ MMLU Verb1s-Top1 Mistral B $0.352\pm 0.004$ $0.132\pm 0.011$ MMLU Verb1s-Top1 Mistral None $0.639\pm 0.006$ $0.255\pm 0.008$ MMLU Verb1s-Top1 Gemma BB (ours) $0.197\pm 0.006$ $\mathbf{0.345\pm 0.035}$ MMLU Verb1s-Top1 Gemma HS-BB (ours) $\mathbf{0.151\pm 0.006}$ $0.3\pm 0.035$ MMLU Verb1s-Top1 Gemma S-BB (ours) $0.161\pm 0.006$ $0.227\pm 0.074$ MMLU Verb1s-Top1 Gemma S-B $0.361\pm 0.007$ $0.133\pm 0.005$ MMLU Verb1s-Top1 Gemma S $0.228\pm 0.005$ $0.137\pm 0.004$ MMLU Verb1s-Top1 Gemma B $0.352\pm 0.004$ $0.144\pm 0.011$ MMLU Verb1s-Top1 Gemma None $0.636\pm 0.007$ $0.258\pm 0.006$ BigBench Ling1S-Top1 Mistral BB (ours) $0.195\pm 0.002$ $0.622\pm 0.009$ BigBench Ling1S-Top1 Mistral HS-BB (ours) $\mathbf{0.138\pm 0.006}$ $\mathbf{0.690\pm 0.031}$ BigBench Ling1S-Top1 Mistral S-BB (ours) $0.157\pm 0.010$ $0.641\pm 0.02$ BigBench Ling1S-Top1 Mistral S-B $0.387\pm 0.006$ $0.523\pm 0.006$ BigBench Ling1S-Top1 Mistral S $0.224\pm 0.003$ $0.52\pm 0.004$ BigBench Ling1S-Top1 Mistral B $0.379\pm 0.003$ $0.612\pm 0.017$ BigBench Ling1S-Top1 Mistral None $0.245\pm 0.004$ $0.684\pm 0.007$ BigBench Ling1S-Top1 Gemma BB (ours) $0.198\pm 0.006$ $0.41\pm 0.012$ BigBench Ling1S-Top1 Gemma HS-BB (ours) $\mathbf{0.142\pm 0.009}$ $\mathbf{0.421\pm 0.021}$ BigBench Ling1S-Top1 Gemma S-BB (ours) $0.193\pm 0.002$ $0.321\pm 0.052$ BigBench Ling1S-Top1 Gemma S-B $0.326\pm 0.002$ $0.302\pm 0.005$ BigBench Ling1S-Top1 Gemma S $0.180\pm 0.006$ $0.298\pm 0.005$ BigBench Ling1S-Top1 Gemma B $0.468\pm 0.001$ $0.293\pm 0.014$ BigBench Ling1S-Top1 Gemma None $0.427\pm 0.006$ $0.412\pm 0.01$
Table 2: Performance of our $\beta$-calibration methods, $\beta$-binning (BB), Pooled scaling-$\beta$-binning (S-BB) and Hierarchical scaling-$\beta$-binning (HS-BB)), compared to the baselines, UMD, (B,  (Gupta and Ramdas, 2021)), Platt Scaling (S, (Platt, 1999)), and Scaling-binning (S-B, (Kumar et al., 2019)), and prompt-based approach None (Tian et al., 2023) . The bold entry in the columns for $\text{CE}(h;\beta)$ and AUAC represent the best performing schemes for that metric. However, note that in some case the confidence intervals overlap, especially between our schemes.
@@ A.1 Limitations of Non-discretized Methods
Estimation of $\mathbb{E}[Y|H]$ when $H=h(Q,A)$ is continuous is a difficult task [Kumar et al., 2019]. The confidence must also be discretized for it to achieve guarantees of marginal calibration [Gupta and Ramdas, 2021]. Prompts used to elicit confidence [Tian et al., 2023, Xiong et al., 2024] are not guaranteed to induce a discretized confidence score (both in the average-case- and $\beta$-calibration cases).
@@
Lastly, estimation of $\text{CE}(h)$ for non-discretized methods may involve binning and can underestimate the error (Proposition 3.3 in [Kumar et al., 2019]). In our case, this is a possibility when comparing scaling against the other approaches.
@@ A.2 Uniform-mass-double-dipping Histogram Binning (UMD)
We adapt UMD (Algorithm 1 from  Gupta and Ramdas [2021]) to our notation in Algorithm 4. UMD takes as input calibration data $\mathcal{D}=\{(h_{i},t_{i})\}_{i\in[N]}$, where $h_{i}$’s are (uncalibrated) confidence scores and $t_{i}$’s are the corresponding target labels. The function order-stats returns ordered confidence scores $(h_{(1)},\ldots,h_{(n)}$ where $h_{(1)}<h_{(2)}<\ldots<h_{(n)}$.
@@ A.3 Distribution-free Guarantees
In Gupta et al. [2020], UMD procedure (Algorithm 4) is assumed to take ground truth $y$ as input ($t=y$). Since in our setting, it is possible to pass a proxy ground truth, we now describe a generalization of conditional calibration guarantees of UMD procedure [Gupta and Ramdas, 2021, Theorem 3].
@@
Consider some $b\in\{0,1,\ldots,B\}$ and denote $l=k_{b-1},u=k_{b}$. By Lemma 2 from [Gupta and Ramdas, 2021], the unordered (denoted by $h_{\{\cdot\}}$) confidence scores $h_{\{l+1\}},h_{\{l+2\}},\ldots,h_{\{u+1\}}$ are i.i.d given ${\mathcal{M}}$, with the same conditional distribution as that of $H$ given ${\mathcal{B}}(H)=b$. Therefore, $\tilde{y}_{\{l+1\}},\tilde{y}_{\{l+2\}},\ldots,\tilde{y}_{\{u-1\}}$ are i.i.d given ${\mathcal{S}}$, with the conditional distribution $\text{Bernoulli}(\mathbb{E}[\tilde{Y}\mid{\mathcal{B}}(H)=b]$.
@@
${\mathbb{P}}\left(\forall p\in\text{range}(g_{s}),\left|\mathbb{E}\left[\tilde{Y}|\beta(q,a)=s,g_{s}(h)=p\right]-p\right|\leq\sqrt{\frac{\log(2B_{s}/\alpha)}{2(\lfloor n_{s}/B_{s}\rfloor-1)}}+\nu\leavevmode\nobreak\ \Bigg{|}\leavevmode\nobreak\ \beta(q,a)=s\right)$
@@
${\mathbb{P}}\left(\forall p\in\text{range}(g_{s}),\left|\mathbb{E}\left[\tilde{Y}|\beta(q,a)=s,g_{s}(h)=p\right]-p\right|\leq\epsilon\leavevmode\nobreak\ \Bigg{|}\leavevmode\nobreak\ \beta(q,a)=s\right)\geq 1-\alpha_{s}.$
@@
${\mathbb{P}}\left(\forall s\in{\mathcal{S}},\forall p\in\text{range}(g_{s}),\left|\mathbb{E}\left[\tilde{Y}|\beta(q,a)=s,g_{s}(h)=p\right]-p\right|\leq\epsilon\leavevmode\nobreak\ \Bigg{|}\leavevmode\nobreak\ \beta(q,a)=s\right)$ $\geq 1-\sum_{s\in{\mathcal{S}}}\alpha_{s}$
@@
To remove the assumption that $h_{i}$’s are distinct in Theorem 3.1, we could use the randomized version of UMD (Algorithm 2 in  Gupta and Ramdas [2021]). This adds an additive factor to $\epsilon$ in Theorem 3.1 but can be made arbitrarily small. For simplicity, we have chosen to provide the guarantees for the non-randomized version of UMD.
@@ B.1 Missing Experimental Details
Details on LM Querying. We set the LM temperature to close to 0 to minimize output stochasticity and set max tokens to be able to process the prompt $m(q)$. We include the prompts in Table 5. The variable EXPRESSION__\__LIST in Ling1S-Top1 prompt is taken from Fagen-Ulmschneider [2015].
@@
Training Details. We perform random splitting 8 times using different seeds and use the repeated measures to form our confidence intervals. We set the number of bins $B$ in the UMD baseline and scaling-binning by deriving hyperparameter search ranges from setting different $\epsilon$s in Theorem 3 in [Gupta and Ramdas, 2021]. We describe how we set the number of points per bin $b$ in Algorithm 1 and Algorithm 3 in Section 5. We have also attempted to ensure that for every $B$ we can find in its hyperparameter tuning range, there is a $b$ in its hyperparameter tuning range, such that $B=\lfloor N/b\rfloor$. This correspondence, coupled with the same target variable to optimize, helps isolate the impact of $\beta$ on post-hoc calibration.
@@
Table 6: (Continuation of Table 2) Performance of our $\beta$-calibration methods.
@@
.      Dataset      Prompt      LLM      Calibrator      $\text{CE}(h;\beta)$      AUAC      SciQ      Ling1S-Top1      Mistral      BB (ours)      $\mathbf{0.165\pm 0.003}$      $\mathbf{0.482\pm 0.024}$      SciQ      Ling1S-Top1      Mistral      HS-BB (ours)      $0.175\pm 0.004$      $0.437\pm 0.028$      SciQ      Ling1S-Top1      Mistral      S-BB (ours)      $0.171\pm 0.006$      $0.426\pm 0.034$      SciQ      Ling1S-Top1      Mistral      S-B      $0.466\pm 0.003$      $0.21\pm 0.007$      SciQ      Ling1S-Top1      Mistral      S      $0.255\pm 0.005$      $0.21\pm 0.005$      SciQ      Ling1S-Top1      Mistral      B      $0.457\pm 0.003$      $0.356\pm 0.025$      SciQ      Ling1S-Top1      Mistral      None      $0.366\pm 0.011$      $0.451\pm 0.009$      SciQ      Ling1S-Top1      Gemma      BB (ours)      $0.158\pm 0.011$      $0.357\pm 0.023$      SciQ      Ling1S-Top1      Gemma      HS-BB (ours)      $0.148\pm 0.011$      $0.363\pm 0.032$      SciQ      Ling1S-Top1      Gemma      S-BB (ours)      $\mathbf{0.146\pm 0.006}$      $0.366\pm 0.085$      SciQ      Ling1S-Top1      Gemma      S-B      $0.485\pm 0.003$      $0.194\pm 0.014$      SciQ      Ling1S-Top1      Gemma      S      $0.211\pm 0.008$      $0.191\pm 0.008$      SciQ      Ling1S-Top1      Gemma      B      $0.486\pm 0.003$      $0.195\pm 0.011$      SciQ      Ling1S-Top1      Gemma      None      $0.446\pm 0.011$      $\mathbf{0.403\pm 0.018}$      SciQ      Verb1S-Top1      Mistral      BB (ours)      $0.262\pm 0.009$      $0.428\pm 0.027$      SciQ      Verb1S-Top1      Mistral      HS-BB (ours)      $0.203\pm 0.006$      $\mathbf{0.474\pm 0.048}$      SciQ      Verb1S-Top1      Mistral      S-BB (ours)      $\mathbf{0.182\pm 0.006}$      $\mathbf{0.474\pm 0.095}$      SciQ      Verb1S-Top1      Mistral      S-B      $0.472\pm 0.003$      $0.237\pm 0.006$      SciQ      Verb1S-Top1      Mistral      S      $0.228\pm 0.005$      $0.27\pm 0.02$      SciQ      Verb1S-Top1      Mistral      B      $0.458\pm 0.002$      $0.377\pm 0.012$      SciQ      Verb1S-Top1      Mistral      None      $0.451\pm 0.013$      $0.451\pm 0.015$      SciQ      Verb1S-Top1      Gemma      BB (ours)      $0.268\pm 0.005$      $0.278\pm 0.021$      SciQ      Verb1S-Top1      Gemma      HS-BB (ours)      $0.194\pm 0.006$      $0.386\pm 0.034$      SciQ      Verb1S-Top1      Gemma      S-BB (ours)      $\mathbf{0.182\pm 0.005}$      $0.279\pm 0.023$      SciQ      Verb1S-Top1      Gemma      S-B      $0.477\pm 0.002$      $0.149\pm 0.013$      SciQ      Verb1S-Top1      Gemma      S      $0.243\pm 0.007$      $0.246\pm 0.008$      SciQ      Verb1S-Top1      Gemma      B      $0.464\pm 0.003$      $0.344\pm 0.021$      SciQ      Verb1S-Top1      Gemma      None      $0.495\pm 0.01$      $\mathbf{0.421\pm 0.01}$      TriviaQA      Ling1S-Top1      Mistral      BB (ours)      $0.26\pm 0.003$      $0.467\pm 0.024$      TriviaQA      Ling1S-Top1      Mistral      HS-BB (ours)      $0.195\pm 0.007$      $0.499\pm 0.014$      TriviaQA      Ling1S-Top1      Mistral      S-BB (ours)      $\mathbf{0.184\pm 0.005}$      $\mathbf{0.541\pm 0.021}$      TriviaQA      Ling1S-Top1      Mistral      S-B      $0.441\pm 0.003$      $0.275\pm 0.006$      TriviaQA      Ling1S-Top1      Mistral      S      $0.266\pm 0.005$      $0.28\pm 0.006$      TriviaQA      Ling1S-Top1      Mistral      B      $0.432\pm 0.003$      $0.441\pm 0.009$      TriviaQA      Ling1S-Top1      Mistral      None      $0.326\pm 0.008$      $0.537\pm 0.01$      TriviaQA      Ling1S-Top1      Gemma      BB (ours)      $0.268\pm 0.007$      $0.310\pm 0.034$      TriviaQA      Ling1S-Top1      Gemma      HS-BB (ours)      $0.188\pm 0.02$      $\mathbf{0.359\pm 0.035}$      TriviaQA      Ling1S-Top1      Gemma      S-BB (ours)      $\mathbf{0.18\pm 0.005}$      $0.351\pm 0.051$      TriviaQA      Ling1S-Top1      Gemma      S-B      $0.235\pm 0.015$      $0.310\pm 0.007$      TriviaQA      Ling1S-Top1      Gemma      S      $0.211\pm 0.012$      $0.086\pm 0.005$      TriviaQA      Ling1S-Top1      Gemma      B      $0.468\pm 0.003$      $0.199\pm 0.019$      TriviaQA      Ling1S-Top1      Gemma      None      $0.486\pm 0.01$      $0.349\pm 0.007$      TriviaQA      Verb1S-Top1      Mistral      BB (ours)      $0.242\pm 0.006$      $\mathbf{0.618\pm 0.016}$      TriviaQA      Verb1S-Top1      Mistral      HS-BB (ours)      $0.191\pm 0.01$      $0.573\pm 0.024$      TriviaQA      Verb1S-Top1      Mistral      S-BB (ours)      $\mathbf{0.18\pm 0.006}$      $0.545\pm 0.029$      TriviaQA      Verb1S-Top1      Mistral      S-B      $0.436\pm 0.005$      $0.324\pm 0.007$      TriviaQA      Verb1S-Top1      Mistral      S      $0.27\pm 0.005$      $0.352\pm 0.014$      TriviaQA      Verb1S-Top1      Mistral      B      $0.409\pm 0.005$      $0.525\pm 0.019$      TriviaQA      Verb1S-Top1      Mistral      None      $0.38\pm 0.01$      $0.582\pm 0.011$      TriviaQA      Verb1S-Top1      Gemma      BB (ours)      $0.249\pm 0.006$      $0.381\pm 0.02$      TriviaQA      Verb1S-Top1      Gemma      HS-BB (ours)      $0.192\pm 0.006$      $\mathbf{0.417\pm 0.034}$      TriviaQA      Verb1S-Top1      Gemma      S-BB (ours)      $\mathbf{0.183\pm 0.008}$      $0.337\pm 0.038$      TriviaQA      Verb1S-Top1      Gemma      S-B      $0.444\pm 0.003$      $0.231\pm 0.012$      TriviaQA      Verb1S-Top1      Gemma      S      $0.258\pm 0.008$      $0.225\pm 0.006$      TriviaQA      Verb1S-Top1      Gemma      B      $0.429\pm 0.004$      $0.37\pm 0.015$      TriviaQA      Verb1S-Top1      Gemma      None      $0.527\pm 0.008$      $\mathbf{0.417\pm 0.008}$
Table 7: (Continuation of Table 2) Performance of our $\beta$-calibration methods.

### v2

@@ 6 Conclusions
We proposed QA-calibration, a new notion of calibration which conditions on groups of QA pairs. We propose two new posthoc calibration schemes for LM-elicited confidence scores. Our algorithms are effective on various QA datasets. For future work, we plan to investigate alternative notions of calibration for groups, such as multi-calibration.
@@
Acknowledgements
P.M was funded in part by the Hybrid Intelligence Center, a 10-year programme funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research with grant number 024.004.022.
@@ A.1 Limitations of Non-discretized Methods
Estimation of $\mathbb{E}[Y|H]$ when $H=h(Q,A)$ is continuous is a difficult task (Kumar et al., 2019). The confidence must also be discretized for it to achieve guarantees of marginal calibration (Gupta & Ramdas, 2021). Prompts used to elicit confidence (Tian et al., 2023; Xiong et al., 2024) are not guaranteed to induce a discretized confidence score (both in the average-case- and QA-calibration cases).
@@
Lastly, estimation of $\text{CE}(h)$ for non-discretized methods may involve binning and can underestimate the error (Proposition 3.3 in (Kumar et al., 2019)). In our case, this is a possibility when comparing scaling against the other approaches.
@@ A.2 Uniform-mass-double-dipping Histogram Binning (UMD)
We adapt UMD (Algorithm 1 from  Gupta & Ramdas (2021)) to our notation in Algorithm 4. UMD takes as input calibration data $\mathcal{D}=\{(h_{i},t_{i})\}_{i\in[N]}$, where $h_{i}$’s are (uncalibrated) confidence scores and $t_{i}$’s are the corresponding target labels. The function order-stats returns ordered confidence scores $(h_{(1)},\ldots,h_{(n)}$ where $h_{(1)}<h_{(2)}<\ldots<h_{(n)}$.
@@ A.3 Distribution-free Guarantees
In Gupta et al. (2020), UMD procedure (Algorithm 4) is assumed to take ground truth $y$ as input ($t=y$). Since in our setting, it is possible to pass a proxy ground truth, we now describe a generalization of conditional calibration guarantees of UMD procedure (Gupta & Ramdas, 2021, Theorem 3).
@@
Consider some $b\in\{0,1,\ldots,B\}$ and denote $l=k_{b-1},u=k_{b}$. By Lemma 2 from (Gupta & Ramdas, 2021), the unordered (denoted by $h_{\{\cdot\}}$) confidence scores $h_{\{l+1\}},h_{\{l+2\}},\ldots,h_{\{u+1\}}$ are i.i.d given ${\mathcal{M}}$, with the same conditional distribution as that of $H$ given ${\mathcal{B}}(H)=b$. Therefore, $\tilde{y}_{\{l+1\}},\tilde{y}_{\{l+2\}},\ldots,\tilde{y}_{\{u-1\}}$ are i.i.d given ${\mathcal{S}}$, with the conditional distribution $\text{Bernoulli}(\mathbb{E}[\tilde{Y}\mid{\mathcal{B}}(H)=b]$.
@@
${\mathbb{P}}\left(\forall p\in\text{range}(g_{s}),\left|\mathbb{E}\left[\tilde{Y}|\beta(q,a)=s,g_{s}(h)=p\right]-p\right|\leq\sqrt{\frac{\log(2B_{s}/\alpha)}{2(\lfloor n_{s}/B_{s}\rfloor-1)}}+\nu~{}\Bigg{|}~{}\beta(q,a)=s\right)$
@@
${\mathbb{P}}\left(\forall p\in\text{range}(g_{s}),\left|\mathbb{E}\left[\tilde{Y}|\beta(q,a)=s,g_{s}(h)=p\right]-p\right|\leq\epsilon~{}\Bigg{|}~{}\beta(q,a)=s\right)\geq 1-\alpha_{s}.$
@@
${\mathbb{P}}\left(\forall s\in{\mathcal{S}},\forall p\in\text{range}(g_{s}),\left|\mathbb{E}\left[\tilde{Y}|\beta(q,a)=s,g_{s}(h)=p\right]-p\right|\leq\epsilon~{}\Bigg{|}~{}\beta(q,a)=s\right)$ $\geq 1-\sum_{s\in{\mathcal{S}}}\alpha_{s}$
@@
To remove the assumption that $h_{i}$’s are distinct in Theorem 3.1, we could use the randomized version of UMD (Algorithm 2 in  Gupta & Ramdas (2021)). This adds an additive factor to $\epsilon$ in Theorem 3.1 but can be made arbitrarily small. For simplicity, we have chosen to provide the guarantees for the non-randomized version of UMD.
@@ B.1 Missing Experimental Details
Details on LM Querying. We set the LM temperature to close to 0 to minimize output stochasticity and set max tokens to be able to process the prompt $m(q)$. We include the prompts in Table 5. The variable EXPRESSION__\__LIST in Ling1S-Top1 prompt is taken from Fagen-Ulmschneider (2015).
@@
Training Details.
We perform random splitting 8 times using different seeds and use the repeated measures to form our confidence intervals. We set the number of bins $B$ in the UMD baseline and scaling-binning by deriving hyperparameter search ranges from setting different $\epsilon$s in Theorem 3 in (Gupta & Ramdas, 2021). We describe how we set the number of points per bin $b$ in Algorithm 1 and Algorithm 3 in Section 5. We have also attempted to ensure that for every $B$ we can find in its hyperparameter tuning range, there is a $b$ in its hyperparameter tuning range, such that $B=\lfloor N/b\rfloor$. This correspondence, coupled with the same target variable to optimize, helps isolate the impact of $\beta$ on posthoc calibration.
@@
Table 6: (Continuation of Table 2) Performance of our QA-calibration methods, QA binning (QAB), Pooled scaling QA binning (S-QAB) and Hierarchical scaling QA binning (HS-QAB)), compared to the baselines, UMD, (B,  (Gupta & Ramdas, 2021)), Platt Scaling (S, (Platt, 1999)), and Scaling-binning (S-B, (Kumar et al., 2019)).
@@
Dataset Prompt LLM Calibrator $\text{CE}(h;\beta)$ AUAC SciQ Ling1S-Top1 Mistral QAB (ours) $\mathbf{0.165\pm 0.003}$ $\mathbf{0.482\pm 0.024}$ SciQ Ling1S-Top1 Mistral HS-QAB (ours) $0.175\pm 0.004$ $0.437\pm 0.028$ SciQ Ling1S-Top1 Mistral S-QAB (ours) $0.171\pm 0.006$ $0.426\pm 0.034$ SciQ Ling1S-Top1 Mistral S-B $0.466\pm 0.003$ $0.21\pm 0.007$ SciQ Ling1S-Top1 Mistral S $0.255\pm 0.005$ $0.21\pm 0.005$ SciQ Ling1S-Top1 Mistral B $0.457\pm 0.003$ $0.356\pm 0.025$ SciQ Ling1S-Top1 Mistral None $0.366\pm 0.011$ $0.451\pm 0.009$ SciQ Ling1S-Top1 Gemma QAB (ours) $0.158\pm 0.011$ $0.357\pm 0.023$ SciQ Ling1S-Top1 Gemma HS-QAB (ours) $0.148\pm 0.011$ $0.363\pm 0.032$ SciQ Ling1S-Top1 Gemma S-QAB (ours) $\mathbf{0.146\pm 0.006}$ $0.366\pm 0.085$ SciQ Ling1S-Top1 Gemma S-B $0.485\pm 0.003$ $0.194\pm 0.014$ SciQ Ling1S-Top1 Gemma S $0.211\pm 0.008$ $0.191\pm 0.008$ SciQ Ling1S-Top1 Gemma B $0.486\pm 0.003$ $0.195\pm 0.011$ SciQ Ling1S-Top1 Gemma None $0.446\pm 0.011$ $\mathbf{0.403\pm 0.018}$ SciQ Verb1S-Top1 Mistral QAB (ours) $0.262\pm 0.009$ $0.428\pm 0.027$ SciQ Verb1S-Top1 Mistral HS-QAB (ours) $0.203\pm 0.006$ $\mathbf{0.474\pm 0.048}$ SciQ Verb1S-Top1 Mistral S-QAB (ours) $\mathbf{0.182\pm 0.006}$ $\mathbf{0.474\pm 0.095}$ SciQ Verb1S-Top1 Mistral S-B $0.472\pm 0.003$ $0.237\pm 0.006$ SciQ Verb1S-Top1 Mistral S $0.228\pm 0.005$ $0.27\pm 0.02$ SciQ Verb1S-Top1 Mistral B $0.458\pm 0.002$ $0.377\pm 0.012$ SciQ Verb1S-Top1 Mistral None $0.451\pm 0.013$ $0.451\pm 0.015$ SciQ Verb1S-Top1 Gemma QAB (ours) $0.268\pm 0.005$ $0.278\pm 0.021$ SciQ Verb1S-Top1 Gemma HS-QAB (ours) $0.194\pm 0.006$ $0.386\pm 0.034$ SciQ Verb1S-Top1 Gemma S-QAB (ours) $\mathbf{0.182\pm 0.005}$ $0.279\pm 0.023$ SciQ Verb1S-Top1 Gemma S-B $0.477\pm 0.002$ $0.149\pm 0.013$ SciQ Verb1S-Top1 Gemma S $0.243\pm 0.007$ $0.246\pm 0.008$ SciQ Verb1S-Top1 Gemma B $0.464\pm 0.003$ $0.344\pm 0.021$ SciQ Verb1S-Top1 Gemma None $0.495\pm 0.01$ $\mathbf{0.421\pm 0.01}$ TriviaQA Ling1S-Top1 Mistral QAB (ours) $0.26\pm 0.003$ $0.467\pm 0.024$ TriviaQA Ling1S-Top1 Mistral HS-QAB (ours) $0.195\pm 0.007$ $0.499\pm 0.014$ TriviaQA Ling1S-Top1 Mistral S-QAB (ours) $\mathbf{0.184\pm 0.005}$ $\mathbf{0.541\pm 0.021}$ TriviaQA Ling1S-Top1 Mistral S-B $0.441\pm 0.003$ $0.275\pm 0.006$ TriviaQA Ling1S-Top1 Mistral S $0.266\pm 0.005$ $0.28\pm 0.006$ TriviaQA Ling1S-Top1 Mistral B $0.432\pm 0.003$ $0.441\pm 0.009$ TriviaQA Ling1S-Top1 Mistral None $0.326\pm 0.008$ $0.537\pm 0.01$ TriviaQA Ling1S-Top1 Gemma QAB (ours) $0.268\pm 0.007$ $0.310\pm 0.034$ TriviaQA Ling1S-Top1 Gemma HS-QAB (ours) $0.188\pm 0.02$ $\mathbf{0.359\pm 0.035}$ TriviaQA Ling1S-Top1 Gemma S-QAB (ours) $\mathbf{0.18\pm 0.005}$ $0.351\pm 0.051$ TriviaQA Ling1S-Top1 Gemma S-B $0.235\pm 0.015$ $0.310\pm 0.007$ TriviaQA Ling1S-Top1 Gemma S $0.211\pm 0.012$ $0.086\pm 0.005$ TriviaQA Ling1S-Top1 Gemma B $0.468\pm 0.003$ $0.199\pm 0.019$ TriviaQA Ling1S-Top1 Gemma None $0.486\pm 0.01$ $0.349\pm 0.007$ TriviaQA Verb1S-Top1 Mistral QAB (ours) $0.242\pm 0.006$ $\mathbf{0.618\pm 0.016}$ TriviaQA Verb1S-Top1 Mistral HS-QAB (ours) $0.191\pm 0.01$ $0.573\pm 0.024$ TriviaQA Verb1S-Top1 Mistral S-QAB (ours) $\mathbf{0.18\pm 0.006}$ $0.545\pm 0.029$ TriviaQA Verb1S-Top1 Mistral S-B $0.436\pm 0.005$ $0.324\pm 0.007$ TriviaQA Verb1S-Top1 Mistral S $0.27\pm 0.005$ $0.352\pm 0.014$ TriviaQA Verb1S-Top1 Mistral B $0.409\pm 0.005$ $0.525\pm 0.019$ TriviaQA Verb1S-Top1 Mistral None $0.38\pm 0.01$ $0.582\pm 0.011$ TriviaQA Verb1S-Top1 Gemma QAB (ours) $0.249\pm 0.006$ $0.381\pm 0.02$ TriviaQA Verb1S-Top1 Gemma HS-QAB (ours) $0.192\pm 0.006$ $\mathbf{0.417\pm 0.034}$ TriviaQA Verb1S-Top1 Gemma S-QAB (ours) $\mathbf{0.183\pm 0.008}$ $0.337\pm 0.038$ TriviaQA Verb1S-Top1 Gemma S-B $0.444\pm 0.003$ $0.231\pm 0.012$ TriviaQA Verb1S-Top1 Gemma S $0.258\pm 0.008$ $0.225\pm 0.006$ TriviaQA Verb1S-Top1 Gemma B $0.429\pm 0.004$ $0.37\pm 0.015$ TriviaQA Verb1S-Top1 Gemma None $0.527\pm 0.008$ $\mathbf{0.417\pm 0.008}$
Table 7: (Continuation of Table 2) Performance of our QA-calibration methods, QA binning (QAB), Pooled scaling QA binning (S-QAB) and Hierarchical scaling QA binning (HS-QAB)), compared to the baselines, UMD, (B,  (Gupta & Ramdas, 2021)), Platt Scaling (S, (Platt, 1999)), and Scaling-binning (S-B, (Kumar et al., 2019)).